OpenStudy (anonymous):

PLEASE HELP!!! The sum of two numbers is 28. The larger number, y, is 4 more than the smaller number, x. Write two equations for this system and solve to find x and y. Give your answer as an ordered pair without parentheses or spaces, as in x,y.

OpenStudy (michele_laino):

From the text of your exercise, we can write: \[x+y=28\] and \[y=4x\]

OpenStudy (anonymous):

wait what?

OpenStudy (campbell_st):

the equations are \[x + y = 28~~~and~~~ y = x+4\]

OpenStudy (michele_laino):

I'm sorry, I was wrong, I think that @campbell_st is right!